Let (N) (a) E(N3N4) (b) E(X3X4) (c) E(S3S4) 20 be a Poisson process with parameter A 2. Find the following:
Answer #1

Pg Noo Answer: - let (Ni b e a poisson process with parameler a) El N3 Ny) con that kle know that NCE) N(4+5)= NE) (NIE)+N (4pague that TE Follcas E[Net). NICZ+5)] = 9++ (atz? +asat 0 In our questions and sal, 2=2. Substitute these values in Egn o E9 No Xi is ith Inter arrival time, which is Independendo [[X3X4) = f(x3). £(X4) - trebuie - 114. Here, f(x): Exponential donbd ELB2,4): F{ - EALA (Žait x)} - & xjp + x4103 f$ + x4 (21+%94%)} =£(80%+ ( *184) + (2224) + f(x334) - Vay( 53) + [6(53))+